OpenStudy (anonymous):

A lander vehicle on a distant planet records the fact that it takes 2.86s for a ball to fall freely 11.66m, starting from rest. What is the acceleration due to gravity on that planet? Express your answer in m/s^2. I am getting -1.425, but according to the hw, the answer is wrong :/.

OpenStudy (anonymous):

S=1/2at^2 11.66m=1/2*a*2.86^2 a=? you can calculate I suppose.

OpenStudy (aravindg):

hi nikvist can u hlp me?